In 2010 the Pandora Box Company made a rights issue at €5 a share of one new share for every four shares held. Before the issue there were 10 million shares outstanding and the share price was €6. Suppose that the company had decided to issue new stock at €4 rather than €5.

  

a.

How many new shares would it have needed to sell to raise the same sum of money? (Enter your answer in whole numbers not in millions.)

  

  Number of shares   

  

b.

The rights issue gives the shareholder the opportunity to buy one new share for less than the market price. What was the value of this opportunity? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  

  Value of the right €   

  

c.

What was the prospective stock price after the issue? (Round your answer to 2 decimal places.)

  

  Prospective stock price €   

  

d.

How far could the total value of the company fall before shareholders would be unwilling to take up their rights? (Enter your answer in millions.)

  

  Total value of the company €   million  
Answer #1

a Right Price = PO = 5 Initial Shares outstanding = NO = 10000000 Rights issued = N1 = (1/4)*NO = 10000000/4 = 2500000 Money

Please give positive vote , let me know if you need any further information regarding the question